OpenStudy (anonymous):

Members of the math team are each given a prize for their participation in a contest. There are 7 calculators, 8 pencil sets, and 5 compasses. The prizes are wrapped in separate, identical boxes. Since Miguel scored highest in the contest, he gets to draw first. What is the probability of Miguel getting a compass? A. 1 over 20 B. 1 over 10 C. 1 over 4 D. 1 over 3 D.

OpenStudy (sepeario):

C - as 5/the total, 20 = 5/20 = 1/4